# State-insensitive wavelengths for light shifts and photon scattering from Zeeman states

### v1

@@ Abstract
Atoms are not two-level systems, and their rich internal structure often leads to complex phenomena in the presence of light. Here, we analyze off-resonant light scattering including the full hyperfine and magnetic structure. We find a set of frequency detunings where the atomic induced dipole is the same irrespective of the magnetic state, and where two-photon transitions that alter the atomic state turn off. For alkali atoms and alkaline-earth ions, if the hyperfine splitting is dominated by the magnetic dipole moment contribution, these detunings approximately coincide. Therefore, at a given “magical” detuning, all magnetic states in a hyperfine manifold behave almost identically, and can be traced out to good approximation. This feature prevents state decoherence due to light scattering, which impacts quantum optics experiments and quantum information applications.
@@ I Introduction
However, the internal electronic structure of atoms (with fine, hyperfine, and magnetic levels) prevents them from behaving as true two-level systems. This often leads to open transitions and complex dynamics involving many levels [15, 16]. While this rich level structure allows for state preparation through optical pumping [17], the realization of scalable quantum information platforms [18, 19, 20, 21, 22, 23, 24] as well as the exploration of spin models with larger than spin-1/2 particles [25, 26, 27, 28, 29, 30, 31, 32], it is in many cases undesired, leading to qubit decoherence and heating. However, restricting atoms to a specific subset of levels is challenging, as it requires initial state preparation and either strict adherence to selection rules to prevent leakage out of a two-level subspace, or continual repumping (e.g., optically) to recover emitters that have left the desired level subspace.
@@
Under certain conditions, however, atoms behave identically irrespective of their internal state. One example is that of state-independent, or “magic”, trapping. Here, the ground and excited states of a transition are both trapped by light of a particular wavelength and polarization such that the induced ac Stark shifts are the same for both levels [33, 34, 35, 36, 37]. This magic wavelength can thus be used to create potentials for both states that have a consistent frequency separation in the presence of amplitude noise, can minimize heating, and thus limit decoherence effects [33, 34, 35]. Atoms held in such potentials enable the realization of higher precision optical atomic clocks and quantum information protocols [38, 35, 39, 40, 41].
@@
In this paper, we find conditions for state-independent photon scattering by multilevel atomic emitters. At a “magic detuning” all magnetic states in a given ground-state hyperfine manifold scatter incoming light in exactly the same manner. Such state-independent scattering requires two criteria to be met: the amplitude and polarization of the induced dipole must be identical for all ground states, and the atom must not change its internal ground state as it scatters. We show that these requirements yield three separate conditions for the polarizability tensor. For a large class of atomic species, these conditions are all met at approximately the same frequency. This coalescence requires scattering on a $D_{2}$ line via three excited states, the hyperfine splitting to be dominated by the nuclear magnetic dipole contribution, and for the atomic spectrum not to feature other relevant lines. We find optimal detunings to achieve state-independent optical responses of alkali atoms and alkaline-earth ions, and find that scattering for the most optimal species, ${}^{133}$s, can be state-independent at up to a part in $10^{7}$. We show this condition also means the ac Stark shift becomes state-independent, and discuss its relevance to cavity QED experiments.
@@
Our paper proceeds as follows. In Sec. II, we introduce our formalism to describe the response of an atom to an external field, the generalized polarizability tensor. In Sec. III, we introduce the conditions needed such that the atomic magnetic state becomes irrelevant. In Sec. IV, we find the detunings that satisfy those introduced conditions. In Sec. V we show that for atoms where the hyperfine structure is dominated by the nuclear magnetic dipole moment, these detunings coalesce to a single “magic detuning”. In Sec. VI we discuss how to optimize the detuning to minimize unwanted scattering, and then perform that optimization for alkali atoms and alkaline-earth ions in Sec. VII. We connect these results to the ac Stark shift in Sec. VIII.1. In Sec. VIII.2, we discuss its relevance to cavity QED experiments and, in particular, to the experiment performed in Ref. [42].
@@
Figure 1: Scattering from a $D_{2}$-line. (a) Atoms are initiated in a hyperfine manifold $F$ and driven (blue) off-resonantly on the $D_{2}$ line with $F^{\prime}\in\{F-1,F,F+1\}$, scattering light (red). Two types of scattering processes that can occur are depicted. The atom can Rayleigh scatter, returning to its original state, or Raman scatter, changing state during the process. (b,c) An atom in an $F=1$ state is driven by a linearly $x$-polarized field and scatters light. (b) For a generic detuning, an atom prepared in a superposition of states becomes entangled with the scattered field, as the polarization and amplitude of the field depend on the magnetic state. (c) At a magic detuning, the generated field is independent of the atomic internal state.
@@ II Generalized polarizability tensor
We consider an atom in a hyperfine state $F$, with $12F+$ 1 magnetic sublevels $m_{F}\in[-F,\,F]$. In the absence of an external magnetic field, these sublevels are degenerate. The atom is driven by a monochromatic field of frequency $\omega$,
$\hat{\mathbf{E}}(\mathbf{r})=\hat{\epsilon}\,\hat{E}_{0}^{(+)}(\mathbf{r})\,\mathrm{e}^{-i\omega t}+\mathrm{c.c},$ (1)
where $\hat{\epsilon}$ is the polarization vector, $\hat{E}_{0}^{(+)}$ is the positive-frequency component of the field, and $\mathbf{r}$ is the position of the atom. The light frequency is close to (but detuned from) a $D_{2}$ line with hyperfine levels $F^{\prime}\in\,\{F-1,\,F,\,F+1\}$, as shown in Fig. 1. The environment is assumed to be frequency selective – e.g. a resonant optical cavity – such that scattering from the $F^{\prime}$ levels to other ground hyperfine states can be neglected.
@@
where $\mu,\nu$ are polarization indices, $d_{\nu(\mu)}$ is the atomic dipole operator projected onto polarization $\nu(\mu)$, and $\Delta_{F^{\prime}}=\omega-\omega_{F^{\prime}}$ is the detuning of the drive field from the $F-F^{\prime}$ resonance transition frequency. The above expression is derived under the rotating wave approximation, which is valid for $\Delta_{F^{\prime}}\ll\omega_{F^{\prime}}$. The polarizability tensor is normalized such that an atom in state |$\ket{F,m_{F}}$ at position $\mathbf{r}$ scatters an input field to an output field at position $\mathbf{r}^{\prime}$ with polarization components
$\hat{E}^{(+)}_{\mathrm{out}}(\mathbf{r}^{\prime})=\mu_{0}\omega^{2}\sum\limits_{\chi,\mu,\nu}G_{\chi\mu}(\mathbf{r}^{\prime},\mathbf{r},\omega)\alpha_{\mu\nu}(F,m_{F},\omega)E^{(+)}_{\nu}(\mathbf{r}).$ (3)
In the above equation, $G_{\chi\mu}(\mathbf{r}^{\prime},\mathbf{r},\omega)$ is the Green’s tensor of the electromagnetic environment evaluated at the atomic position.
This standard form of the polarizability tensor does not describe all possible scattering processes, as it does not allow for the atom to change internal state. This can be justified by assuming scattering only due to stimulated emission by the drive and that there is a large Zeeman splitting making such Raman transitions strongly off resonant [43]. However, in the absence of a large magnetic field, one should instead consider a more general map that includes different initial and final magnetic sublevels. Inspired by Eq. (2), we define the “generalized polarizability tensor” as
@@
where we have dropped the subscript $F$ from the ground atomic states, simply using $m$ for the sake of clarity.
In general, the state of the atom becomes correlated with the field, and the positive-frequency component of the field operator is related to the atomic coherence operators between Zeeman sublevels as
$\hat{\mathbf{E}}^{(+)}_{\mathrm{out}}(\mathbf{r}^{\prime})=\mu_{0}\omega^{2}\sum\limits_{\chi,\nu,\mu,m,n}G_{\chi\mu}$ $(\mathbf{r}^{\prime},\mathbf{r},\omega)\,\alpha^{nm}_{\mu\nu}(F,\omega)$
×$\times\hat{E}^{(+)}_{\nu}(\mathbf{r})\ket{F,m}\bra{F,n}.$ | . (5)
Elements of the generalized polarizability tensor that are diagonal in the atomic state (i.e., $n=m$), correspond to Rayleigh scattering processes and are simply given by the polarizability tensor of Eq. (2). Off-diagonal elements (i.e., $n\neq m$) correspond to Raman processes. Both these processes can lead to correlations between the atom and scattered field, and so (even in the absence of Raman processes), an atom initially in a superposition of $m$ states will decohere as correlations develop. For the field to be independent of |$\ket{m}\bra{n}$ |, we require the generalized polarizability tensor to be of the form
@@
i.e., for $\alpha^{nm}_{\mu\nu}$ to be diagonal in the magnetic state (this is, only non-zero when $n=m$), and $\alpha^{mm}_{\mu\nu}$ to be independent of $m$ such that the above sum over the atomic states becomes the identity. In the following, we use these criteria to find conditions on the dipole matrix elements, and find detunings at which the atomic state does not become correlated with the field.
Figure 2: Scattering processes that depend on the internal atomic state. (a) The ellipticity of the field may not be preserved in the induced dipole. Due to symmetry, states $m$ and $-m$ will respond in opposite ways, and thus this process is $m$ dependent. (b) The amplitude of the induced dipole may be different for different states. (c) Scattering processes that change the magnetic state can occur. Since not all states can undergo these processes, they are $m$ dependent.
II.1 Choice of polarization basis
The coordinate system can be arbitrarily chosen without loss of generality, and so we choose both the quantization axis of the atoms and the propagation direction of the drive to be $z$. Therefore, the polarization vector of the drive field lies in the $x$-$y$ plane  [^1] Note that this choice of polarization is not the same as that in Ref. [42], where it was assumed the (linear-polarized) drive was $\pi$-polarized. Here, we choose a different polarization axis because it is more convenient for the algebra to work in the $x-y$ plane for arbitrary elliptical polarization. Of course, the physics is unaffected by this choice, and our results will hold for any polarization axis choice.. Within this plane, it is convenient to work in a basis defined by the circular polarization unit vectors $\hat{\epsilon}_{\pm}=(\hat{\epsilon}_{x}\pm{\rm i}\hat{\epsilon}_{y})/\sqrt{2}$. Using this convention, we define a unit polarization vector parallel to that of the incoming light, and one perpendicular, of the form
@@ II.1 Choice of polarization basis
II.2 Evaluating dipole matrix elements
@@
These selections rules can be used to eliminate the sums over $m_{F^{\prime}}$ in the generalized polarizability tensor.
@@ III.1 Rayleigh scattering
The first criteria we require is that the atom Rayleigh-scatters (i.e., $m=n$) at a rate that is independent of the atomic state. For scattering into the parallel mode, we need $\alpha^{mm}_{\parallel\parallel}(F,\omega)=\alpha_{\parallel\parallel}(F,\omega)~{}\forall~{}m$. These terms are given by
@@
For linear polarization, $\cos\theta=\sin\theta$, and for the above expression to be state independent there is only one condition: that the sum of the two different-handed terms is independent of $m$. This condition enforces that the amplitude of the induced dipole must be the same for all states.
@@
Therefore, the difference between the two different-handed terms is zero for the $m=0$ state. This means that for this difference to be independent of $m$, it must be zero for all $m$. This condition enforces that the induced dipole preserves the ellipticity of the input field.

### v2

@@ Abstract
Atoms are not two-level systems, and their rich internal structure often leads to complex phenomena in the presence of light. Here, we analyze off-resonant light scattering including the full hyperfine and magnetic structure. We find a set of frequency detunings where the induced atomic dipole is the same irrespective of the Zeeman state, and where two-photon transitions that alter the atomic state turn off. For alkali atoms and alkaline-earth ions, if the hyperfine splitting is dominated by the magnetic dipole moment contribution, these detunings approximately coincide. Therefore, at a given “magical” detuning, all Zeeman states in a hyperfine manifold behave almost identically, and can be traced out to good approximation. This feature prevents state decoherence due to light scattering, which impacts quantum optics experiments and quantum information applications.
@@ I Introduction
However, the internal structure of atoms – with fine, hyperfine, and Zeeman levels – prevents them from behaving as true two-level systems, instead displaying open transitions and complex dynamics involving many levels [15, 16, 17]. This rich level structure allows for state preparation through optical pumping [18], the realization of scalable quantum information platforms [19, 20, 21, 22, 23, 24, 25] and the exploration of spin models with larger-than-spin-1/2 particles [26, 27, 28, 29, 30, 31, 32, 33], among many applications. Nevertheless, open transitions are in many cases undesired, as they lead to qubit decoherence and heating. However, restricting atoms to a specific subset of levels is challenging, as it requires initial state preparation and either strict adherence to selection rules to prevent leakage out of a two-level subspace, or continual repumping (e.g., optically) to recover atoms that have left the desired level subspace.
@@
Under certain conditions, however, atoms behave identically irrespective of their internal state. One example is that of state-independent, or “magic”, trapping. Here, the ground and excited states of a transition are both trapped by light of a particular wavelength and polarization such that the induced ac Stark shifts are the same for both levels [34]. Magic wavelengths create potentials for both states that have a consistent frequency separation in the presence of amplitude noise, therefore minimizing heating and limiting decoherence. Atoms held in magic wavelength potentials enable the realization of higher precision optical atomic clocks and more efficient quantum information protocols [35, 36, 37, 38, 39]. Magic wavelengths were first found for clock transitions in strontium [40, 35, 41] and since extended to other atomic clock candidates such as ytterbium [42, 43] and mercury [44]. Magic wavelengths have also been found for optical transitions in cesium [45] and in alkaline-earth ions [46, 47, 48]. Moreover, magic potentials generated by multiple frequencies have been implemented for optical transitions in rubidium [49, 50, 51].
@@
In this paper, we extend the concept of magic wavelengths to Zeeman states in the same ground-state hyperfine manifold. We find conditions for state-independent photon scattering, showing that at a particular “magic detuning” all Zeeman states scatter incoming light in exactly the same manner. Such state-independent scattering requires two criteria to be met: the amplitude and polarization of the induced dipole must be identical for all Zeeman states, and the atom must not change its Zeeman state as it scatters. We show that these requirements yield three separate conditions for the polarizability tensor. For a large class of atomic species, these conditions are all met at approximately the same frequency. This coalescence requires: (1) scattering via three excited states, (2) for the hyperfine splitting to be dominated by the nuclear magnetic dipole contribution, and (3) for the atomic spectrum not to feature other relevant lines. We find optimal detunings to achieve state-independent optical responses of alkali atoms and alkaline-earth ions, and find that scattering for the most optimal species, 133Cs, can be state-independent at up to a part in $10^{7}$. We show this condition also implies that the ac Stark shift becomes state-independent, and discuss its relevance to cavity quantum electrodynamics (QED) experiments, where this effect was recently demonstrated and utilized [52].
@@
Our paper proceeds as follows. In Sec. II, we introduce our formalism to describe the response of an atom to an external field, the generalized polarizability tensor. In Sec. III, we introduce the conditions needed such that the atomic Zeeman state becomes irrelevant. In Sec. IV, we find the detunings that satisfy those conditions. In Sec. V we show that, for atoms where the hyperfine structure is dominated by the nuclear magnetic dipole moment, these detunings coalesce to a single “magic detuning”. In Sec. VI we discuss how to optimize the detuning to minimize unwanted scattering, and then perform that optimization for alkali atoms and alkaline-earth ions in Sec. VII. We connect these results to the ac Stark shift in Sec. VIII.1. In Sec. VIII.2, we discuss its relevance to cavity QED experiments and, in particular, to the experiment performed in Ref. [52].
@@
Figure 1: Scattering from a hyperfine-split line. (a) The atom is initially in a (ground state) hyperfine manifold $F$ (shown here as $F=1$ for simplicity, but generically $F\geq 1$). The atom is then driven (blue) off-resonantly via the three lines with $F^{\prime}\in\{F-1,F,F+1\}$, scattering light (red). Two types of scattering processes that can occur are depicted. The atom can Rayleigh scatter, returning to its original state, or Raman scatter, changing state during the process. (b,c) An atom in an $F=1$ state is driven by a linearly $x$-polarized field and scatters light. (b) For a generic detuning, an atom prepared in a superposition of states becomes entangled with the scattered field, as the polarization and amplitude of the field depend on the Zeeman state. (c) At a magic detuning, the generated field is independent of the Zeeman state.
@@ II Generalized polarizability tensor
We consider an atom driven by the electric field of monochromatic light. In the following, we describe the mathematical structure used in the rest of the paper. We discuss the level structure of the atom, how to characterize its response to the field, the coordinate basis used to perform the characterization, and how to evaluate the necessary dipole matrix elements.
II.1 Atomic structure
The ground state angular momentum of an atom is defined by its nuclear spin $I$ and its (ground state) total electron angular momentum $J$. Assuming $I+J\geq 1$, the ground state is split into different hyperfine manifolds labeled by $|I-J|\leq F\leq I+J$. In the absence of a magnetic field, each of these ground hyperfine manifolds has a $12F+$ 1 degeneracy due to its Zeeman sublevels. We consider light scattering via an excited state with total electron angular momentum $J^{\prime}$. This state is also subject to hyperfine splitting, such that scattering is via three excited hyperfine manifolds $F^{\prime}\in\,\{F-1,\,F,\,F+1\}$. This structure can be found in transitions such that $J^{\prime}=J+1\,\,\forall\,F$, or such that $J^{\prime}=J$ and $F\leq I+J-1$ or $J^{\prime}=J-1$ and $F\leq I+J-2$. The necessary structure is most readily found in alkali atoms and singly ionized alkaline-earth ions scattering from the $D_{2}$ line with $J^{\prime}=J+1$. The simplest configuration of an $F=1$ hyperfine manifold is shown in Fig. 1(a). The environment is assumed to be frequency selective – e.g. a resonant optical cavity – such that scattering from the $F^{\prime}$ levels to other ground hyperfine manifolds can be neglected.
II.2 Atomic response to an electric field
We consider that the atom is driven by a monochromatic field of frequency $\omega$. The input field operator reads
$\hat{\mathbf{E}}_{\mathrm{in}}(\mathbf{r})=\hat{\mathbf{E}}_{\mathrm{in}}^{(+)}(\mathbf{r})\,\mathrm{e}^{-i\omega t}+\mathrm{c.c},$ (1)
where $\hat{\mathbf{E}}_{\mathrm{in}}^{(+)}$ is the positive-frequency component of the field with polarization $\hat{\epsilon}$, and $\mathbf{r}$ is the position of the atom. Here, for generality, the field is considered as an operator, but the treatment described in the rest of the manuscript is valid and unchanged if applied to a classical amplitude.
@@ II.2 Atomic response to an electric field
where $\mu,\nu$ are polarization indices, $d_{\nu(\mu)}$ is the atomic dipole operator projected onto polarization $\nu(\mu)$, and $\Delta_{F^{\prime}}=\omega-\omega_{F^{\prime}}$ is the detuning of the drive field from the $F-F^{\prime}$ resonance transition frequency. The above expression is derived under the rotating wave approximation, valid for $\Delta_{F^{\prime}}\ll\omega_{F^{\prime}}$.
The polarizability tensor encodes the dipole induced by the field on the atom and, therefore, also the field scattered by the atom. Specifically, an atom in state |$\ket{F,m_{F}}$ at position $\mathbf{r}$ subject to the input field defined in Eq. (1) produces an output field at position $\mathbf{r}^{\prime}$ with $\chi$-polarized positive frequency component
$\hat{\text{E}}^{(+)}_{\mathrm{out},\chi}(\mathbf{r}^{\prime})=\mu_{0}\omega^{2}\sum\limits_{\mu,\nu}G_{\chi\mu}(\mathbf{r}^{\prime},\mathbf{r},\omega)\alpha_{\mu\nu}(F,m_{F},\omega)\hat{\text{E}}^{(+)}_{\mathrm{in},\nu}(\mathbf{r}).$ (3)
In the above equation, $\mu_{0}$ is the permeability of free space, $G_{\chi\mu}(\mathbf{r}^{\prime},\mathbf{r},\omega)$ is the Green’s tensor of the electromagnetic environment that describes the propagation of the output field from the atomic position to some position $\mathbf{r}^{\prime}$  [^1] As the Green’s tensor describes propagation of the electromagnetic field, its precise spectral and spatial form depends on the boundary conditions. For instance, the spatial dependence of the Green’s tensor for the guided mode (of wavevector $k$) of a waveguide reads $G(z^{\prime},z,\omega)\propto\exp({{\rm i}k|z^{\prime}-z|})$, while for a Fabry-Pérot cavity it reads $G(z^{\prime},z,\omega)\propto\cos(kz)\cos(kz^{\prime})$ [78], where in both cases we have assumed light propagation only along the $z$-axis., and $\hat{E}^{(+)}_{\mathrm{in},\nu}(\mathbf{r})$ is the $\nu$-polarized positive frequency component of the input field.
The standard form of the polarizability tensor does not describe all possible scattering processes, as it does not allow for the atom to change from its initial Zeeman state. The absence of state-changing transitions is usually justified by assuming they are strongly off resonant due to a large Zeeman splitting [54]. However, in the absence of a large magnetic field, one should instead consider a more general map that includes different initial and final Zeeman state. Inspired by Eq. (2), we define the “generalized polarizability tensor” as
@@
where we have dropped the subscript $F$ from the ground Zeeman states, simply using $m$ and $n$ for the sake of clarity.
In general, the state of the atom becomes correlated with the field, and the positive-frequency component of the field operator is related to the atomic coherence operators between Zeeman states as
$\hat{\text{E}}^{(+)}_{\mathrm{out},\chi}(\mathbf{r}^{\prime})=\mu_{0}\omega^{2}\sum\limits_{\nu,\mu,m,n}G_{\chi\mu}$ $(\mathbf{r}^{\prime},\mathbf{r},\omega)\,\alpha^{nm}_{\mu\nu}(F,\omega)$
×$\times\hat{\text{E}}^{(+)}_{\mathrm{in},\nu}(\mathbf{r})\ket{F,m}\bra{F,n}.$ | . (5)
Elements of the generalized polarizability tensor that are diagonal in the atomic state (i.e., $n=m$), correspond to Rayleigh scattering processes and are simply given by the polarizability tensor of Eq. (2). Off-diagonal elements (i.e., $n\neq m$) correspond to Raman processes. Both these processes can generate correlations between the atom and scattered field, and so (even in the absence of Raman processes), an atom initially in a superposition of $m$ states will decohere as correlations develop. For the field to be independent of |$\ket{m}\bra{n}$ |, we require the generalized polarizability tensor to be of the form
@@
i.e., for $\alpha^{nm}_{\mu\nu}$ to be diagonal in the Zeeman state (this is, only non-zero when $n=m$), and to be independent of $m$, i.e., $\alpha^{mm}_{\mu\nu}(F,\omega)=\alpha_{\mu\nu}(F,\omega)~{}\forall~{}m$ such that the above sum over the atomic states becomes the identity. Note that this condition means that for all Zeeman states we recover the conventional polarizability tensor of Eq. (2), with the additional constraint that it is independent of $m$. Thus, if a state-insensitive wavelength can be found, the polarizability tensor can be calculated by choosing any $m$ value in Eq. (2). In the following, we use these criteria to find conditions on the dipole matrix elements, and find detunings at which the atomic state does not become correlated with the field.
Figure 2: Scattering processes that depend on the Zeeman state. (a) The ellipticity of the field may not be preserved in the induced dipole. Due to symmetry, states $m$ and $-m$ will respond in opposite ways, and thus this process is $m$ dependent. (b) The amplitude of the induced dipole may be different for different states. (c) Scattering processes that change the Zeeman state can occur. Since not all states can undergo these processes, they are $m$ dependent.
II.3 Choice of polarization basis
The coordinate system can be arbitrarily chosen without loss of generality, and so we choose both the quantization axis of the atoms and the propagation direction of the drive to be $z$. Therefore, the polarization vector of the drive field lies in the $x$-$y$ plane  [^2] Note that this choice of polarization is not the same as that in Ref. [52], where it was assumed the (linear-polarized) drive was $\pi$-polarized. Here, we choose a different polarization axis because it is more convenient for the algebra to work in the $x-y$ plane for arbitrary elliptical polarization. Of course, the physics is unaffected by this choice, and our results will hold for any choice of polarization axis.. Within this plane, it is convenient to work in a basis defined by the circular polarization unit vectors $\hat{\epsilon}_{\pm}=(\hat{\epsilon}_{x}\pm{\rm i}\hat{\epsilon}_{y})/\sqrt{2}$. Using this convention, we define a unit polarization vector parallel to that of the incoming light, and one perpendicular, of the form
@@ II.3 Choice of polarization basis
II.4 Evaluating dipole matrix elements
@@
These selections rules can be used to eliminate the sums over $m_{F^{\prime}}$ in the generalized polarizability tensor of Eq. II.2.
@@ III.1 Rayleigh scattering
The first required condition is that the atom Rayleigh-scatters (i.e., $m=n$) at a rate that is independent of its Zeeman state. For scattering into the parallel mode, we need $\alpha^{mm}_{\parallel\parallel}(F,\omega)=\alpha_{\parallel\parallel}(F,\omega)~{}\forall~{}m$. These terms are given by
@@
For linear polarization, $\cos\theta=\sin\theta$, and for the above expression to be state independent there is only one condition: that the sum of the two different-handed terms is independent of $m$. This condition restricts the amplitude of the induced dipole to be identical for all states.
@@
Therefore, the difference between the two different-handed terms is zero for the $m=0$ state. For this difference to be independent of $m$, it must be zero for all $m$. This condition enforces that the induced dipole preserves the ellipticity of the input field.